Find the equations of the parabola satisfying the given conditions. The vertex of each is at the origin.
1. Focus (0, 0.4)
2. Focus (2.5, 0)

Answer by lwsshak3(11628)

Find the equations of the parabola satisfying the given conditions. The vertex of each is at the origin.
1. Focus (0, 0.4)
Parabola opens upward
axis of symmetry: x=0 or y-axis
basic form of equation: x^2=4py
p=0.4(distance from vertex to focus on the axis of symmetry)
4p=1.6
equation: x^2=1.6y
..
2. Focus (2.5, 0)
Parabola opens rightward
axis of symmetry: y=0 or x-axis
basic form of equation: y^2=4px
p=2.5(distance from vertex to focus on the axis of symmetry)
4p=10
equation: y^2=10y